Find the general solution of the differential equation: dy = 2t y +(cost)et2


1
Expert's answer
2021-08-30T16:28:42-0400

"\frac{dy}{dt} = 2t y +(cost)e^{t^2}\\
\frac{dy}{dt} -2t y =(cost)e^{t^2}\\
\text{This is a linear differential equation.}\\
\text{Integrating factor, IF}=e^{\int-2tdt}=e^{-t^2}\\
\text{Therefore, solution is given by}\\
ye^{-t^2}=\int e^{-t^2}(cost)e^{t^2}dt\\
ye^{-t^2}=\int cost \space dt\\
ye^{-t^2}=sint+c\\
y=e^{t^2}(sint+c)"